Consider the following hypothesis test:


H₀: p ≥ 0.75 

Hα: p<0.75



A sample of 400 items was selected. Compute the p-value and state your conclusion for each of the following sample results. Use α=.05.

Round your answers to four decimal places.

a. p̅=0.69


b. p̅=0.72


c. p̅=0.7


d. p̅=0.79
Answer #1

This is the left tailed test .

The null and alternative hypothesis is

H0 : p = 0.75

Ha : p < 0.75

n =400

a ) \hat p = 0.61

P0 = 0.75

1 - P0 = 1 - 0.75 = 0.25

Test statistic = z

= \hat p - P0 / [\sqrtP0 * (1 - P0 ) / n]

= 0.61 - 0.75/ [\sqrt0.75 * 0.25 / 400]

= −6.466

Test statistic = z = −6.47

P-value = 0

\alpha = 0.05

P-value < \alpha

0 < 0.05

Reject the null hypothesis .

There is sufficient evidence to suggest that < 0.75

b ) \hat p = 0.72

P0 = 0.75

1 - P0 = 1 - 0.75 = 0.25

Test statistic = z

= \hat p - P0 / [\sqrtP0 * (1 - P0 ) / n]

= 0.72 - 0.75/ [\sqrt0.75 * 0.25 / 400]

= −1.386

Test statistic = z = −1.39

P-value = 0.0829

\alpha = 0.05

P-value ≥\alpha

0.0829 ≥ 0.05

Do not reject the null hypothesis .

There is sufficient evidence to suggest that < 0.75

c ) \hat p = 0.70

P0 = 0.75

1 - P0 = 1 - 0.75 = 0.25

Test statistic = z

= \hat p - P0 / [\sqrtP0 * (1 - P0 ) / n]

= 0.70 - 0.75/ [\sqrt0.75 * 0.25 / 400]

= −2.309

Test statistic = z = −2.31

P-value = 0.0105

\alpha = 0.05

P-value < \alpha

0.0105< 0.05

Reject the null hypothesis .

There is sufficient evidence to suggest that < 0.75

d ) \hat p = 0.79

P0 = 0.75

1 - P0 = 1 - 0.75 = 0.25

Test statistic = z

= \hat p - P0 / [\sqrtP0 * (1 - P0 ) / n]

= 0.79 - 0.75/ [\sqrt0.75 * 0.25 / 400]

= 1.848

Test statistic = z = 1.85

P-value = 0.9677

\alpha = 0.05

P-value ≥ \alpha

0.9677 ≥0.05

Do not reject the null hypothesis

There is sufficient evidence to suggest that < 0.75

thank you for the question..........kindly rate...........it helps me a lot